# a Benchmark Dataset for Diagnoses and Deterioration Prediction in Emergency Medicine

### v3

@@ Abstract
Background: A clinically meaningful comparative assessment of medical decision support in emergency care is challenging due to a lack of appropriate datasets with multimodal input modalities and comprehensive prediction task. This hampers measurable progress in the field.
Results: We introduce a dataset based on MIMIC-IV, a benchmarking protocol, and initial results for evaluating multimodal decision support in the emergency department (ED). We use diverse data modalities from the first 1.5 hours after patient arrival, including demographics, biometrics, vital signs, lab values, and electrocardiogram waveforms. We analyze 1443 clinical labels across two contexts: predicting diagnoses and patient deterioration. Our diagnostic model achieves an AUROC score over 0.8 in a statistically significant manner for 609 out of 1428 conditions, including cardiac conditions like myocardial infarction and non-cardiac conditions such as renal disease and diabetes. The deterioration model scores above 0.8 in a statistically significant manner for 14 out of 15 targets, including critical events like cardiac arrest, mechanical ventilation, intensive care unit admission, as well as short- and long-term mortality. Furthermore, we provide one of the first robust demonstrations of the significant impact of raw waveform input data on model performance.
Conclusions: This study highlights the proposed dataset as a unique resource to foster progress towards measurable progress in the domain of algorithmic decision support in emergency care. The presented multimodal baseline models showcase the potential of diagnostic decision support in the field and provide strong incentives for including raw waveform data.
@@
Medical Decision Support; Multimodal Data; Emergency Department; Diagnostic Prediction; Patient Deterioration.
journal: general
{keypoints*}
  • Introduction of a novel dataset based on MIMIC-IV for benchmarking multimodal decision support in the emergency department.
  • Evaluation of 1443 clinical labels for diagnostic and patient deterioration predictions, achieving high AUROC scores.
  • Demonstration of significant performance improvements by incorporating raw ECG waveform data.
  • Highlighting the dataset’s comprehensive feature set collected within the first 1.5 hours after patient arrival.
  • Potential impact on improving decision-making processes in acute and emergency medicine.
@@ 1 Introduction
Artificial intelligence (AI) is an emerging field that has significantly enhanced healthcare and medicine [1] in areas such as precision medicine [2] and drug discovery [3]. The implementation of AI models is particularly relevant in acute and emergency medicine, where clinicians must address critical conditions within short time frames to make optimal clinical decisions. AI models can improve acute care by enabling early diagnosis [4], proposing tailored diagnostic workups [5], predicting admissions (including ward or intensive care unit (ICU)) [6], estimating survival rates [7], and providing faster and low-cost diagnoses [8, 9]. There has been exponential growth in the number of scientific publications on new AI models. However, many of these studies have limitations, such as predicting only a limited set of specific conditions [10, 4], offering short time horizons for survival predictions, and requiring costly diagnostic tests [11].
@@
Healthcare datasets are essential for advancing medical research and innovation. However, much of this valuable data remains undisclosed due to commercial interests and privacy concerns. Despite these hurdles, there has been a recent surge in publicly accessible datasets aimed at facilitating the development and validation of machine learning models to tackle complex healthcare challenges [12, 13, 14, 15, 16]. Nonetheless, these datasets still have limitations, including size constraints [13], lack of open-source availability [17, 18], and narrow scopes in terms of prediction tasks [13, 14, 18, 15, 17, 16]. Benchmarking is crucial for evaluating the performance and reliability of novel AI-models, ensuring their applicability and robustness in real-world clinical settings. In this study, we investigate multimodal decision support for emergency department (MDS-ED) using an open-source biomedical multimodal dataset which comprises various medical feature modalities collected within the first 1.5 hours of patient arrival at the ED department such as demographics, biometrics, vital parameters trends, laboratory trends, and electrocardiogram (ECG) waveforms, to predict a diverse and large range of clinical tasks beyond traditional predictive scenarios, such as cardiac and non-cardiac diagnoses, as well as patient deterioration such as mortalities, ICU admission, organ-specific failure, and mechanical ventilation to name a few.
2 Analyses
@@ 2.1 Benchmarking predictive performance
Table 1 presents the predictive performance of both unimodal and multimodal models across the diagnoses and deterioration settings. For diagnoses prediction, the multimodal model that integrates ECG waveforms and clinical routine data achieved a macro average across all areas under the respective receiver operating curves (AUROC) (macro-AUROC) of 0.8256 (0.8222,0.8288), which is notably higher than the rest of the models. Similarly, in the deterioration prediction task, the multimodal model that integrates ECG waveforms and clinical routine data achieved an AUROC of 0.9115 (0.8991,0.9222), which also overcomes the rest of the models. In the following discussion, we restrict ourselves to the discussion of the most comprehensive multimodal model based on ECG waveforms and clinical routine data.
@@ 2.2 Task-dependent predictive performance
Table 2 demonstrates the diagnoses that can be predicted most accurately by the model, which we organized based on ICD chapters. The algorithm’s predictive performance ranges from an AUROC of 0.7405 for the musculoskeletal system and connective tissue (XIII) chapter to 0.8761 for the circulatory system (IX) chapter. In total, the model predicted 609 out of 1,428 individual ICD diagnoses with high accuracy, which we define in this work as conditions where the lower bound of the 95% bootstrap confidence interval for AUROC exceeds 0.80.
@@
Table 3 showcases the deterioration model’s predictive performance across clinical deterioration, ICU admission, as well as both short and long-term mortality with many AUROC scores significantly exceeding the 0.80 threshold. Overall, for clinical deterioration, the model achieves an AUROC of 0.9070, whereas, for ICU admissions, the model reports an overall AUROC of 0.9063, and for mortality predictions also exhibits high performance, with an overall AUROC of 0.9168.
@@ 3.1 Impact of data modalities
From the results presented in the previous section, we can draw several conclusions: Firstly, the results demonstrate that multimodal models, which integrate diverse data types, offer superior performance in both diagnostic and deterioration tasks (row 4&5 vs. the rest). Secondly, in the diagnoses task as well as in the deterioration task, the use of ECG raw waveforms instead of ECG features improves the performance in a statistically significant manner (row 4 vs. row 5), finding which is not in line with [14]. To the best of our knowledge, this is the first statistically robust demonstration of the added value of raw ECG waveform input against ECG features for clinically relevant prediction tasks such as diagnoses and deterioration prediction. Thirdly, for the unimodal models, in the deterioration task, the clinical routine data model outperforms ECG features only and ECG waveforms only, however, in the diagnoses task, the ECG waveforms only outperforms the other 2 settings, we hypothesize that for the deterioration task, the clinical routine data apart of including a rich set of clinical features (demographics, biometrics, vital parameters trends, and laboratory values trends) against only an single ECG either in features or waveform, it also includes trends over time which aligns with the task definition of deterioration. Despite this, we believe that a single ECG snapshot can achieve high performances for both tasks, but also we believe that the inclusion of multiple ECGs over time instead of just a single snapshot would allow us to capture more meaningful deterioration and potentially diagnoses trends.

### v4

# Enhancing clinical decision support with physiological waveforms – a multimodal benchmark in emergency care

@@ Abstract
Background: AI-driven prediction algorithms have the potential to enhance emergency medicine by enabling rapid and accurate decision-making regarding patient status and potential deterioration. However, the integration of multimodal data, including raw waveform signals, remains underexplored in clinical decision support.
Methods: We present a dataset and benchmarking protocol designed to advance multimodal decision support in emergency care. Our models utilize demographics, biometrics, vital signs, laboratory values, and electrocardiogram (ECG) waveforms as inputs to predict both discharge diagnoses and patient deterioration.
Results: The diagnostic model achieves area under the receiver operating curve (AUROC) scores above 0.8 for 609 out of 1,428 conditions, covering both cardiac (e.g., myocardial infarction) and non-cardiac (e.g., renal disease, diabetes) diagnoses. The deterioration model attains AUROC scores above 0.8 for 14 out of 15 targets, accurately predicting critical events such as cardiac arrest, mechanical ventilation, ICU admission, and mortality.
Conclusions: Our study highlights the positive impact of incorporating raw waveform data into decision support models, improving predictive performance. By introducing a unique, publicly available dataset and baseline models, we provide a foundation for measurable progress in AI-driven decision support for emergency care.
@@
Deep-learning , Emergency Department , Machine learning , Medical Decision Support , Multimodal Data , Patient Diagnostics , Patient Deterioration.
journal: Computers in Biology and Medicine
url]https://uol.de/en/ai4health
[…]
[label1]organization=AI4Health Division, Carl von Ossietzky Universität Oldenburg, addressline=Ammerländer Heerstraße 114-118, city=Oldenburg, postcode=26129, state=Lower Saxony, country=Germany
[…]
[label2]organization=Department of Internal Medicine, Department of Acute Care, and Department of Clinical Pharmacy & Pharmacology, University Medical Center Groningen, addressline=Hanzeplein 1, city=Groningen, postcode=9713, state=Groningen, country=Netherlands
@@ 1 Introduction
Artificial intelligence (AI) is an emerging field that has significantly enhanced healthcare and medicine [1] in areas such as precision medicine [2] and drug discovery [3]. The implementation of AI models is particularly relevant in acute and emergency care, where clinicians must address critical conditions within short time frames to make optimal clinical decisions. AI models can improve acute care by enabling early diagnosis [4], proposing tailored diagnostic workups [5], predicting admissions (including ward or intensive care unit (ICU)) [6], estimating survival rates [7], and providing faster and low-cost diagnoses [8, 9]. There has been exponential growth in the number of scientific publications on new AI models. However, many of these studies have limitations, such as predicting only a limited set of specific conditions [10, 4], offering short time horizons for survival predictions, and requiring costly diagnostic tests [11].
Despite their potential, current state-of-the-art AI models for forecasting specific medical conditions rely primarily on standard clinical features (e.g., demographics, laboratory results) and neglect raw physiological signals such as waveforms. These signals are known to be highly informative for a diverse set of diseases [12] and represent one example of raw sensor data. They hold the promise of providing complementary information to standard clinical features to enhance the accuracy and robustness of prediction models. The prediction models used vary widely, as noted in a recent meta-review [13] in the context of sepsis prediction, but typically cannot be compared directly due to differences in experimental setup. This highlights the need for structured benchmarking protocols. Unlike mere evaluation, which often focuses on fixed metrics for a single study, benchmarking establishes common datasets, tasks, and performance baselines, ensuring that advancements in AI-driven clinical decision support are measurable, reproducible, and comparable across studies. Benchmarking, as one component of standardized performance evaluation, also plays a crucial role in validation and certification processes. These processes represent a significant hurdle that decision support systems must overcome to reach clinical practice.
@@
Healthcare datasets are essential for advancing medical research and innovation. However, much of this valuable data remains undisclosed due to commercial interests and privacy concerns. Despite these hurdles, there has been a recent surge in publicly accessible datasets aimed at facilitating the development and validation of machine learning models to tackle complex healthcare challenges [12, 14, 15, 16, 17]. Nonetheless, these datasets still have limitations, including size constraints [14], lack of open-source availability [18, 19], and narrow scopes in terms of prediction tasks [14, 15, 19, 16, 18, 17].
In this study, we investigate a multimodal decision support for emergency department (MDS-ED) using an open-source biomedical multimodal dataset which comprises various medical feature modalities collected within the first 1.5 hours of patient arrival at the ED department such as demographics, biometrics, vital parameters trends, laboratory trends, and electrocardiogram (ECG) waveforms, to predict a diverse and large range of clinical tasks beyond traditional predictive scenarios, such as cardiac and non-cardiac diagnoses, as well as patient deterioration such as mortalities, ICU admission, organ-specific failure, and mechanical ventilation to name a few. Compared to approaches using a single data modality, which may offer a limited view of the patient’s condition, we demonstrate that fusing multiple data sources is critical for improving decision-making accuracy in the complex environment of emergency departments. This multimodal approach is essential for predicting immediate as well as long-term outcomes, enabling more comprehensive and accurate assessments of patient status. To summarize, we put forward the following technical contributions:
  1. 1.
    We propose the MDS-ED dataset along with a corresponding benchmarking protocol which stands out from literature approaches as detailed in Section 2.2.
  2. 2.
    We put forward strong unimodular baseline models based on best-practices from the literature, see Section 2.3.
  3. 3.
    We propose powerful multimodal prediction models that outperform these strong unimodal baselines in all the scenarios considered, as demonstrated in Section 3. To our knowledge, this is the first robust demonstration of the added value of incorporating raw waveforms into clinical prediction models in an acute care setting.
2 Methods
2.1 Clinical workflow and dataset creation
Figure 1: Pipeline outlines the MDS-ED clinical workflow, which involves feature collection encompassing patient demographics, biometrics (such as height, weight, and BMI), vital parameters and trends, laboratory values and trends, and ECG waveform data to address two clinically relevant prediction scenarios: predicting patient discharge diagnoses out of 1428 cardiac and non-cardiac ICD10-CM codes and predicting patient deterioration according to 15 clinical deterioration measures.
Figure 1 presents a schematic illustration of the proposed MDS-ED pipeline which task focuses on a specific workflow with high clinical relevance, where for each ED visit, it encompasses a set of features collected from a window of 90 minutes from the patient’s arrival at the ED to predict patient discharge diagnoses and deterioration through the stay. MDS-ED was created by linking ECG waveforms MIMIC database, specifically, MIMIC-IV-ECG [20] dataset to clinical features and outcomes as clinical ground truth from the MIMIC-IV and MIMIC-IV-ED dataset [21]. The MIMIC dataset was chosen for its extensive collection of patient records across diverse demographic groups, encompassing thousands of features from clinical notes, lab results, vital signs, and full waveform data. This multimodal breadth enhances AI models predictive performance and clinical decision support.
Figure 2: Schematic representation summarizing the creation process of the MDS-ED dataset [22] underlying this work. The process starts from four different source datasets (MIMIC-IV-ECG, MIMIC-IV, MIMIC-IV-ED, and MIMIC-IV-ECG-ICD) from which we select patients aged 18 years or older where an ECG was collected within the first 90 minutes of arrival at the ED department. Different target values collected at different intervals. On the resulting samples, we procedded to apply a features outlier removal which is primarily error-based, excluding unrealistic values, never-registered extremes, or negative values when the minimum is zero, see Appendix C and [22] for details. Similarly, an imputation mask creation in which we apply median imputation from the train set to validation and test, adding binary masks to indicate imputed values, helping the model learn missingness patterns. Finally, we capture trends through engineered features such as summary statistics (mean, median, min, max, standard deviation), first and last values, rate of change, and slope of fitted linear model based on values within the first 90 minutes after arrival. The resulting MDS-ED dataset comprises 1428 diagnostic labels and 15 deteriorations labels across 129,057 samples from 71,098 patients collected from 121,195 unique visits. The input features cover a single 10s, 12-lead ECG in addition to 470 tabular features (excluding binary masking columns).
The MDS-ED dataset was constructed by integrating multiple sources to create a comprehensive resource for emergency medicine research. This process involved merging ECG recordings with diverse patient characteristics and clinical outcomes. Specifically, diagnostic ICD-10 codes and stratified splits were obtained from MIMIC-IV-ECG-ICD [12], while ECG recordings were sourced from MIMIC-IV-ECG [20]. Clinical data were extracted from MIMIC-IV [21], utilizing tables related to patient admissions, diagnoses, laboratory tests, ICU stays, and procedures. Additionally, emergency department-specific information, including stays, diagnoses, medication administration, vital signs, and triage details, was gathered from MIMIC-IV-ED [21]. The dataset creation process in visualized schematically in Figure 2. A complete description of the MDS-ED dataset along with the full preprocessing code for its creation is available on PhysioNet[22].
Prediction tasks and targets
In this work, we broadly categorize all prediction tasks into two groups: patient discharge diagnoses and patient deterioration. For discharge diagnoses, we follow MIMIC-IV-ECG-ICD [12], which proposed a framework to predict discharge diagnoses based on a single electrocardiogram (ECG). To this end, we frame the task as a multilabel classification task in terms of international classification of diseases clinical modification 10 (ICD-10 CM) codes. As in [12], we convert ICD-9 to ICD-10 CM codes where necessary, convert codes to 5 digits, and propagate codes to parent codes up to the third digit. In this way, we obtain a total set of 1428 significantly populated ICD-10 CM codes covering a wide range of cardiac and non-cardiac conditions. For patient deterioration, we build on a consensus definition [23] and investigate 15 deterioration events in a multilabel setting after the patient arrived at different intervals and after the 1.5-hour feature collection window: Six subtasks cover clinical deterioration within 24 hours and cover severe hypoxemia, extracorporeal membrane oxygenation (ECMO), use of vasopressors, use of inotropes, mechanical ventilation, and in-hospital cardiac arrest (IHCA). Two subtasks cover ICU admission within 24 hours and or within the entire stay, and finally, 7 subtasks cover mortality prediction at different horizons ranging from inpatient mortality, 24 hours, 7 days, 28 days, 90 days, 180 days, to 365 days.
While ordinary predictive models often focus on immediate deterioration, our approach extends beyond standard scenarios by considering a broader temporal scope and a more comprehensive set of clinical events. Long-term prediction tasks, such as mortality at extended time horizons, pose unique challenges due to the evolving nature of patient conditions, while discharge diagnoses rely on accumulated clinical knowledge throughout the hospital stay and could be highly influential if identified early. By addressing both immediate and long-term risks, our approach ensures a more holistic and clinically relevant decision support framework. With the two main groups of diagnoses and deterioration, we aim to cover a wide spectrum of patient monitoring in the ED as well as long-term mortality across 1,443 unique target labels. See Table 7 in the supplementary material for a detailed discussion of the prediction targets.
Features
Train-test splits
MDS-ED consists of data from 71,098 patients, encompassing 121,195 unique visits, and a total of 129,057 samples. We leverage the stratified splits provided along with the dataset accompanying [12], which includes stratification based on gender, age bins, and discharge diagnoses. We distribute the total number of 20 stratified folds into train, validation, and test following ratios of 18:1:1.
2.2 Related work
Group Detail MIMIC-ED-Assist VitalML MEME EHRSHOT MC-BEC MDS-ED
Source [14] [15] [19] [16] [18] This work
Population ED ED Longitudinal Longitudinal ED ED
Size Patients 25714 N/A 947028 6739 63389 71098
Visits 32356 19847 N/A 921499 102731 121195
Features Demographics
Biometrics
Vital parameters ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$) ✓($\mathbb{T}$)
Lab. values ✓($\mathbb{T}$) ✓($\mathbb{T}$)
Waveforms ✓($\mathbb{E}$)
Chief complaint
Medications
Tasks Labels 3 3 4 15 7 1443
Availability Open source
  • 1.
Table 1 contains a direct comparison with related works in terms of population, dataset size, features, availability, and number of target labels. [14] proposed the MIMIC-ED-Assist benchmark to reduce ED length of stay by flagging high-risk patients using triage features. [15] proposed VitalML to predict deterioration in the next 90 minutes from the first 15 minutes of monitoring. [19] proposed a multiple embedding model for EHR data (MEME) to generate pseudo-notes from tabular data and predict ED disposition, discharge location, ICU admission, and inpatient mortality utilizing a natural language processing approach. [16] proposed a few shot foundational model EHRSHOT approach to predict operational outcomes, developed diagnoses, as well as anticipate lab test results and image findings. [18] proposed MC-BEC to predict patient deterioration, disposition, and revisit from features collected including diagnoses from the first 15 minutes after the patient is assigned to a room.
2.3 Model architectures
To propose an initial benchmark, we experiment with diverse settings based on data modalities input such as (1) clinical routine data, (2) ECG features only, (3) ECG waveforms only, (4) ECG features + clinical routine data, and (4) ECG waveforms + clinical routine data. For waveforms (3) and (5), inspired by recent successful applications in the field of physiological time series [12, 25, 26] we employed an S4 classifier based on structured state space models with four layers [27] (relying on hyperparameter choices from prior work[26], i.e, using state dimension 8 and model dimension 512), which has been shown to outperform more conventional deep learning architectures, such as modern convolutional neural networks, recurrent neural networks or transformer models, as signal encoder [12, 25, 26]. For the multimodal models in (5), we use a 3-layer MLP as tabular encoder, where the three categorical features (gender, race, acuity score) are first processed through appropriate embedding layers. The outputs of the signal encoder and the tabular encoder are then fused through concatenation. We apply median imputation using training set statistics. In addition, we introduce a binary column per input variable to indicate imputed values. These binary columns are provided as additional inputs next to the original tabular features. In A in the supplementary material, we present an ablation study comparing different imputation strategies. A systematic exploration of other tabular encoder models or more complex multimodal fusion schemes beyond the basic choices from above is deferred to future work. For (1), (2) and (4) we fit extreme gradient boosting (XGBoost) decision trees, which widely accepted state-of-the-art for prediction models on tabular data. We used standard hyperparameter defaults commonly applied in gradient-boosting methods. Additional details on model architectures and model hyperparameters can be found in D Table 11.
2.4 Training and evaluation
For the S4 models, we used AdamW [28] as the optimizer, with both learning rate and weight decay set to 0.001, and maintained a constant learning rate schedule. The training was conducted with a batch size of 64 samples over 20 epochs The hyperparameter settings coincide with those from an earlier unimodal investigation [12], which underscores the robustness of our findings without excessive hyperparameter tuning. Overfitting was mitigated through model selection based on validation performance, making the total number of epochs irrelevant as long as it was sufficiently large for convergence. The training objective was to minimize binary cross-entropy loss. For all the models, we evaluate performances on the macro average across all areas under the respective receiver operating curves (AUROC) (macro AUROC). To assess statistical uncertainty resulting from the test set’s finite size and specific composition, we employ empirical bootstrap on the test set with $n=1000$ iterations. We report 95% confidence intervals for both macro AUROC and individual label AUROCs. Given [12] evaluation protocol, we use only the first record per visit per patient to avoid bias. We use macro AUROC as the primary metric due to its status as most widely used ranking-based metric. It characterize a model’s overall discriminative power without the necessity to fix custom decision thresholds. The suitability of AUROC compared to the other commonly used metric in the presence of label imbalance, the area under the precision recall curve (AUPRC), was confirmed in a recent theoretical and empirical work [29]. Additional details of training and evaluation procedures can be found in D.
3 Results
@@ 3.1 Benchmarking predictive performance
Table 2 presents the predictive performance of both unimodal and multimodal models across the diagnoses and deterioration settings. For diagnoses prediction, the multimodal model that integrates ECG waveforms and clinical routine data achieved a macro average across all areas under the respective receiver operating curves (AUROC) (macro-AUROC) of 0.8256 (0.8222,0.8288), which is notably higher than the rest of the models. Similarly, in the deterioration prediction task, the multimodal model that integrates ECG waveforms and clinical routine data achieved an AUROC of 0.9115 (0.8991,0.9222), which also overcomes the rest of the models. In the following discussion, we restrict ourselves to the discussion of the most comprehensive multimodal model based on ECG waveforms and clinical routine data. In B, we present a more finegrained comparison across different ICD10 chapters and specific deterioration targets, where the multimodal deep learning consistently outperforms unimodal baseline models.
@@ 3.2 Task-dependent predictive performance
Table 3 demonstrates the diagnoses that can be predicted most accurately by the model, which we organized based on ICD chapters. The algorithm’s predictive performance ranges from an AUROC of 0.7405 for the musculoskeletal system and connective tissue (XIII) chapter to 0.8761 for the circulatory system (IX) chapter. In total, the model predicted 609 out of 1,428 individual ICD diagnoses with high accuracy, which we define in this work as conditions where the lower bound of the 95% bootstrap confidence interval for AUROC exceeds 0.80. A more fine-grained comparison of model performance across different input settings in the deterioration prediction task reveals notable trends in Table 5, where the inclusion of ECG waveforms alongside clinical routine data consistently improves performance over ECG features and clinical routine data, with relative improvements ranging from 2.36% to 13.06%. The highest gains are observed for XII (Skin, 13.06%), VI (Nervous, 10.75%), and XIX (Injury, Poisoning, 9.95%), suggesting that continuous waveform signals provide critical insights over features into these non-cardiac groups. In contrast, smaller improvements are seen for III (Blood, 2.36%) and IV (Endocrine, 5.25%), indicating that static ECG features might already capture most of the predictive signal in these cases.
@@
Table 4 showcases the deterioration model’s predictive performance across clinical deterioration, ICU admission, as well as both short and long-term mortality with many AUROC scores significantly exceeding the 0.80 threshold. Overall, for clinical deterioration, the model achieves an AUROC of 0.9070, whereas, for ICU admissions, the model reports an overall AUROC of 0.9063, and for mortality predictions also exhibits high performance, with an overall AUROC of 0.9168. In Tables 5 and 6 in the supplementary material, we present a more fine-grained comparison of model performance across different input settings for the deterioration prediction task. As in the diagnostic case, the multimodal model leveraging waveforms outperforms the corresponding model leveraging only ECG features in addition to clinical routine data in all cases, most notably in the clinical deterioration category with a relative improvement of more than 9%. On the most finegrained level in Table 6, one observes relative improvements ranging from 1.41% (365-day mortality) to 38.08% (severe hypoxemia). The most substantial gains occur for severe hypoxemia (38.08%) and inotrope administration (16.57%), suggesting that waveform-based inputs contribute significantly to an early identification of respiratory failure and cardiac dysfunction. Conversely, for longer-term mortality outcomes (e.g., 90-day to 365-day), the added benefit of waveform data diminishes, as patient trajectories become increasingly influenced by non-acute factors that are already reliably captured by clinical routine features and ECG features.
@@ 4.1 Impact of data modalities
From the results presented in the previous section, we can draw several conclusions: Firstly, the results demonstrate that multimodal models, which integrate diverse data types, offer superior performance in both diagnostic and deterioration tasks (row 4&5 vs. the rest). Secondly, in the diagnoses task as well as in the deterioration task, the use of ECG raw waveforms instead of ECG features improves the performance in a statistically significant manner (row 4 vs. row 5), finding which is not in line with [15]. To the best of our knowledge, this is the first statistically robust demonstration of the added value of raw ECG waveform input against ECG features for clinically relevant prediction tasks such as diagnoses and deterioration prediction. Thirdly, for the unimodal models, in the deterioration task, the clinical routine data model outperforms ECG features only and ECG waveforms only, however, in the diagnoses task, the ECG waveforms only outperforms the other two settings, we hypothesize that for the deterioration task, the clinical routine data apart of including a rich set of clinical features (demographics, biometrics, vital parameters trends, and laboratory values trends) against only an single ECG either in features or waveform, it also includes trends over time which aligns with the task definition of deterioration. Despite this, we believe that a single ECG snapshot can achieve high performances for both tasks, but also we believe that the inclusion of multiple ECGs over time instead of just a single snapshot would allow us to capture more meaningful deterioration and potentially diagnoses trends. Fourthly, direct comparison to literature results is challenging due to the lack of publicly available datasets and the corresponding scarcity of models that integrate both clinical routine data and waveform data. Among existing studies, [18] is one of the few incorporating waveform-based models, but differences in dataset construction and task formulation limit direct comparisons. Consequently, from a methodological perspective, we evaluate our approach using state-of-the-art models, including an ensemble tree-based model for tabular data and a signal encoder with proven efficacy, as previously presented in [12].